Hi I'm doing a test revision on some problems I got wrong on an assessment 3 weeks ago. This is one of the four problems I got wrong:
4z-2(z+5)<=8z
(<= stands for less than or equal to)
Thanks for taking the time to help me!

Found 2 solutions by Fombitz, mananth:
Answer by Fombitz(32388)
Answer by mananth(16946)

4z-2(z+5)<=8z
4z-2z-10<=8z
..
-8z -8z
4z-8z-2z-10<=0
..
-6z-10<=0
+10 +10
-6z-10+10<=10
-6z<=10
/-6
z>=10/-6
z>=-5/3